OpenStudy (anonymous):

HeLp 911 What is the similarity ratio of two cubes with volumes of 64 cubic centimeters and 125 cubic centimeters? 2:3 4:5 1:5 16:25

OpenStudy (anonymous):

my answr is b

OpenStudy (phi):

if you know the volumes of two similar figures, you know that their lengths are similar. the ratio of their lengths is the cubed root of the ratio of their volumes \[\sqrt[3]{\frac{64}{125}}\] the cubed root of 64 is 4: 4*4*4= 16*4= 64 the cubed root of 125 is 5: 5*5*5= 25*5= 125 so the answer is 4/5